Hi, I am really confused with my homework and my teacher won't help me so... My quesions is about of course Proportions and Dimensional Analysis. I sort of understand the proportions, but not the other one. Three of my problems are:
a.) 50 meters per second to kilometers per hour.
b.) 0.025 day to seconds. (Now this one I really don't understand)
c.) 1200 ounces to tons (16 oz=1 lb; 2000 lb=1 ton).
And here is one problem for a proportion question:
0.926 meter/1 yard = 20 meters/x yards.

Answer by rfer(16322)

a)
50m per sec = km per hr
360 sec per hr
50*360=18000 m per hr=18km per hr
b)
what part of a day is 2.5% in seconds
24 hrs per day times 360 seconds per hr=8640sec per day
8640*.025=216 seconds
c)
1200/16=75 lbs
75/2000=.0375 tons
d)
0.926/1=20/x
cross multiply
0.926x=20
x=21.6 yds